OpenStudy (anonymous):

two fair dice are rolled repeatedly until the total number of spots that comes up is divisible by 3. what is the probability that this first happens on a roll number divisble by 3

OpenStudy (anonymous):

probably have to sum a geometric series

OpenStudy (anonymous):

we can do this for sure i think that you have to sum a series, but it is not hard

OpenStudy (anonymous):

P(sum divisible by 3) = 1/3 do you mean i have to do this? (1/3)^3+(1/3)^6+......

OpenStudy (anonymous):

if it first happens on the third roll, probability is \[\frac{2}{3}\times \frac{2}{3}\times \frac{1}{3}=\frac{4}{27}\]

OpenStudy (anonymous):

since the probability it is divisible by 3 is \(\frac{1}{3}\) and the probability it is not divisible by 3 is \(\frac{2}{3}\) so that particular sequence represents "not , not, is"

OpenStudy (anonymous):

then the sequence for getting it for the first time on the sixth roll is \[\frac{2}{3}\times \frac{2}{3}\times \frac{2}{3}\times\frac{2}{3}\times\frac{2}{3}\times\frac{1}{3}\]

OpenStudy (anonymous):

and so on the series will look like \[\frac{4}{27}+\frac{4}{27}\times \frac{8}{27}+\frac{4}{27}\times \left(\frac{8}{27}\right)^2+...\]

OpenStudy (anonymous):

you can add this up via \[\frac{a}{1-r}\] with \(a=\frac{4}{27}\) and \(r=\frac{8}{27}\)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i get \(\frac{4}{19}\) does that seem right?

OpenStudy (anonymous):

yes. that's what I got too

OpenStudy (anonymous):

ok